# Jiazhen Liu, Peihan Li, Yuwei Wu, Gaurav S. Sukhatme, Vijay Kumar, Lifeng Zhou

### v2

@@ Abstract
Multi-robot target tracking finds extensive applications in different scenarios, such as environmental surveillance and wildfire management, which require the robustness of the practical deployment of multi-robot systems in uncertain and dangerous environments. Traditional approaches often focus on the performance of tracking accuracy with no modeling and assumption of the environments, neglecting potential environmental hazards which result in system failures in real-world deployments. To address this challenge, we investigate multi-robot target tracking in the adversarial environment considering sensing and communication attacks with uncertainty. We design specific strategies to avoid different danger zones and proposed a multi-agent tracking framework under the perilous environment. We approximate the probabilistic constraints and formulate practical optimization strategies to address computational challenges efficiently. We evaluate the performance of our proposed methods in simulations to demonstrate the ability of robots to adjust their risk-aware behaviors under different levels of environmental uncertainty and risk confidence. The proposed method is further validated via real-world robot experiments where a team of drones successfully track dynamic ground robots while being risk-aware of the sensing and/or communication danger zones.
@@ 1 Introduction
Deployment of multiple robots to track real targets in practical scenarios remains challenging, as the environment or targets themselves can be dangerous or adversarial. Robots may be subject to attacks resulting in sensor damage, communication interruption, or other system faults. Therefore, tracking performance is no longer the only factor to consider, as the robots must also pay special attention to secure themselves from attacks. [9, 10, 11, 12, 13] have closely investigated multi-robot target tracking in the adversarial setting. [10] designed a distributed strategy based on divide-and-conquer that can protect robots from denial-of-service (DoS) attacks. However, it considers attacks in the worst case and overestimates the number of attacks. [10] then improved the overly conservative strategy with an algorithm that can infer the number of DoS attacks using information from 3-hop neighbors. [9] similarly focused on worst-case attacks but showed through simulations that the algorithm still gives superior performance even when robots are under non-worst-case attacks (such as random attacks). [12] extended the notion of target tracking and studied the resilient information-gathering problem. It proposed an algorithm that sustains resiliency against any number of failures while the robots keep only minimal communication between them.
Though awareness of risk has been investigated in the context of multi-robot systems in these prior works, the worst-case assumption is often adopted, leading to excessively conservative decision-making. In this paper, we alternatively consider risk in a probabilistic sense and model the safety requirement for robots, i.e., the risk level must not exceed a threshold, using chance-based constraints. Our problem is relevant to motion/trajectory planning under uncertainty [14, 15, 16, 17, 18], which are highly relevant to our problem. [14] models an autonomous vehicle as a linear Gaussian system and requires the probability of the vehicle either colliding with obstacles or leaving stay zones to be upper-bounded. First, planning for the vehicle is formulated as a chance-constrained optimization problem, which is, in general, intractable due to its chance-based nature. Then, the problem is reformulated into a disjunctive convex optimization that can be solved to global optimality.[15] presented a probabilistic collision avoidance method to navigate micro aerial vehicles through dynamic obstacles. Linear approximation of the probabilistic collision constraints is applied such that the algorithm runs in real time. Inspired by these works, we apply the approximation techniques to alleviate the computational hardness of evaluating the risk constraints. Our contributions can be summarized as :
@@ 2.2 Sensing danger zone
where pdf is the probability density function, $\mathbf{w}_{i,\mathcal{S}_{l}}=\mathbf{x}_{\mathcal{S}_{l}}-\mathbf{x}_{i}$, and it also follows a Gaussian distribution $\mathcal{N}(\mathbf{\mu}_{\mathcal{S}_{l}}~{}-~{}\mathbf{x}_{i},~{}\mathbf{\Sigma}_{\mathcal{S}_{l}})$. The probability is an integral of a multivariate Gaussian random variable over a disk-shaped region $\{\mathbf{w}|\lVert\mathbf{w}_{i,\mathcal{S}_{l}}\rVert\leq r_{l}\}$, which lacks a closed form solution and thus cannot be directly computed.
@@ 2.3 Communication danger zone
Figure 1: (a) Geometrical illustration for a communication danger zone where the position of the central attacking source is non-deterministic. The uncertainty is reflected as blue radial gradient; (b) Shows the original integration region corresponding to the chance constraints; (c) Linearizes the integration region to compute an upper-bound of the multivariate integral. The line $\mathbf{a}^{\top}\mathbf{x}=b$ is tangent to the circle.
Next, we show the geometric conditions under which the jamming attack will be successful, given that exact positions of the jamming sources are non-deterministic. We use robot $i$, its teammate $j$ for $j\neq i$, and the jamming source in zone $\mathcal{C}_{k}$ as an example, but the idea applies to all triplets. As shown in Fig. 1, when robot $i$ is inside a communication danger zone $\mathcal{C}_{k}$, the jamming source $k$ can attack the communication channel between robot $i$ and its teammate $j$. Let the distance between robot $i$ and $j$ be $c$, and the distance between robot $i$ and the jamming source $k$ be $a$. We encode the degree to which robot $i$ and $j$ can maintain reliable communication using the ratio of distances [19, 20], defined as:
@@ 2.4 Target Tracking Optimization with Danger Zones
With two types of danger zones introduced, we now present a nonlinear optimization program in Eq. 8 to model online target tracking under the threats of sensing and communication attacks. Eq. 8 optimizes tracking performance in the objective function and incorporates safety constraints originating from the danger zones. The optimization over control inputs $\mathbf{u}_{t}=[\mathbf{u}_{1,t},\dots,\mathbf{u}_{M,t}]^{\intercal}$ for the whole robot team is as follows,
@@
The objective function balances between minimizing tracking error (first term) and control efforts (second term) with constant weights $w_{1},w_{2}$. $J(\cdot,\cdot)$ represents the uncertainty in target state estimation. It is a nonlinear function dependent on robot positions at time $t+1$ and predicted target positions at time $t+1$, i.e. $\mathbf{x}_{t+1}$ and $\hat{\mathbf{z}}_{t+1}$. Eq. 8b describes the dynamics constraint of all the robots. $\mathbf{x}_{t+1}=[\mathbf{x}_{1,t},\dots,\mathbf{x}_{M,t}]^{\intercal}$ is the state of all robots at step $t+1$. We adopt a simple linear dynamics model, with process matrix $\mathbf{\Phi}$ and control matrix $\mathbf{\Lambda}$, but switching to more complex ones is straightforward. Eq. 8c is the constraint preventing robots from entering sensing danger zones. Eq. 8d is the constraint to preserve effective inter-robot communication in communication danger zones. We use $a_{ik}$ to denote the distance between robot $i$ and the jamming source of communication danger zone $\mathcal{C}_{k}$, and $c_{i}^{*}$ to denote the $c^{*}$ value of robot $i$ (see Eq. 6), both calculated using robot $i$’s position at time step $t+1$, i.e., $\mathbf{x}_{i,t+1}$. This problem is challenging to solve directly with probabilistic constraints (8c)(8d) and the objective function involved the evaluation towards robots’ position $\mathbf{x}_{t+1}$ and predicted positions of targets $\hat{\mathbf{z}}_{t+1}$. In the following section, we will discuss the formulation and approximation of the optimization problem.
@@ 3.1 Chanced-based Constraints Approximation
$\text{Prob}(\mathbf{a}^{\top}\mathbf{x}\leq b)\leq\delta\iff\mathbf{a}^{\top}\mathbf{\mu}-b\geq c,$ (9)
where $c=\text{erf}^{-1}(1-2\delta)\sqrt{2\mathbf{a}^{\top}\mathbf{\Sigma}\mathbf{a}}$, $\text{erf}(\cdot)$ is the standard error function.
The chance constraints in Eq. 8c and Eq. 8d can be converted into convex deterministic constraints, that are convex and consequently simplify the computation process. The integration areas for both types of danger zones are disk-shaped, as illustrated in Fig. 1. We employ a linear approximation of the integration area to calculate a conservative upper bound of the probability. This approximation is demonstrated in Fig. 1, where the integration boundary is approximated by a line $\mathbf{a}^{\top}\mathbf{x}=b$ tangent to the disk.
@@
In other words, it requires the approximate probability to be bounded by $\epsilon_{1}$. $\mathbf{a}_{i,\mathcal{S}_{l}}=(\mathbf{\mu}_{\mathcal{S}_{l}}-\mathbf{x}_{i})/\lVert\mathbf{\mu}_{\mathcal{S}_{l}}-\mathbf{x}_{i}\rVert$ is a unit vector. The chance constraint in Eq. 10 has the same form as Eq. 9, equivalent to the subsequent deterministic constraint,
@@ 3.2 Mission Objective
The objective function Eq. 8a involves minimizing the uncertainty of target state estimation when robots are at resultant position $\mathbf{x}_{t+1}$. We let $f(\cdot,\cdot)$ be the trace of the estimation covariance matrix for target positions calculated using $\mathbf{x}_{t+1}$ and $\mathbf{z}_{t+1}$ with the latter being the positions of targets at time $t+1$. In other words, the first term in Eq. 8a tries to find the next-step robots’ positions that minimize the tracking error and control effort utilizing the team’s estimated information of the targets. In the upcoming part, we show that the trace value is non-convex with respect to the optimization variables.
To estimate the targets’ positions recursively, we assume each robot is equipped with one range and one bearing sensor, which are nonlinear in the positions of targets. We use $\mathbf{z}_{t}$ to denote the positions of targets, with the position of a particular target $j$ being $\mathbf{z}_{j,t}=[x_{T_{j,t}},~{}y_{T_{j,t}}]^{\top}$, and use $\mathbf{x}_{i,t}=[x_{i,t},~{}y_{i,t}]^{\top}$ to denote the position of robot $i$. Based on the derivation in [21], robot $i$’s range and bearing measurements of target $j$ are
$d_{ij}=\sqrt{\Delta x_{ij}^{2}+\Delta y_{ij}^{2}}\qquad\theta_{ij}=\text{arctan}(\frac{\Delta y_{ij}}{\Delta x_{ij}})-\phi_{i},$ (14)
where $\Delta x_{ij}=x_{T_{j}}-x_{i},{}\Delta y_{ij}=y_{T_{j}}-y_{i}$, and $\phi_{i}$ is the orientation of robot $i$. We drop the time step $t$ here for simplicity of notation. Robot $i$’s measurement of target $j$ is thus
@@
Here, $n_{d_{ij}}$ and $n_{\theta_{ij}}$ are noises in range and bearing measurements. We assume both follow zero-mean Gaussian distributions. Specifically, we let $n_{d_{ij}}\sim\mathcal{N}(0,R_{d,ij})$. The measurement noise covariance, $R_{d,ij}$, is a nonlinear function of the distance between robot $i$ and target $j$. We define it through its inverse as
$R^{-1}_{d,ij}=a_{d}\cdot\text{exp}(-\lambda_{d}\lVert\mathbf{x}_{i}-\mathbf{z}_{j}\rVert),$ (16)
where $a_{d}$ and $\lambda_{d}$ are constant parameters determining the characteristic of the noise. The bearing measurement noise covariance, $R_{\theta,ij}$, is defined similarly. The team-level measurement noise covariance matrix with respect to target $j$, i.e., $\mathbf{R}_{j}$, is thereby constructed as $\mathbf{R}_{j}^{-1}=\text{diag}(R^{-1}_{d,1j},R^{-1}_{\theta,1j},R^{-1}_{d,2j},R^{-1}_{\theta,2j},\cdots,R^{-1}_{d,Mj},R^{-1}_{\theta,Mj})$, and it follows that $\mathbf{R}=\text{diag}(R_{1},R_{2},\cdots,R_{N})$ represents the noise covariance matrix of the whole robot team measuring all targets.
To compute the uncertainty of target position estimation, we need the measurement matrix for a combination of range and bearing sensors. By [21], we derive robot $i$’s measurement matrix of target $j$ as
@@
where $\mathbf{h}_{d_{ij}}$ and $\mathbf{h}_{\theta_{ij}}$ are column vectors corresponding to the range and bearing measurements. We define $\mathbf{p}_{ij}=\mathbf{x}_{i}-\mathbf{z}_{j}$, then the two measurement vectors are
@@
$\mathbf{J}$ is the $2\times 2$ rotational matrix applied at $-\frac{\pi}{2}$. We can concatenate the measurement matrices of target $j$ from each robot vertically to obtain the whole team’s measurement matrix,
$\mathbf{H}_{j}=[\mathbf{H}_{1j};\mathbf{H}_{2j};\cdots;\mathbf{H}_{Mj}].$ (19)
The robot team’s measurements of all targets are thereby constructed as $\mathbf{H}=\text{diag}(\mathbf{H}_{1},\mathbf{H}_{2},\cdots,\mathbf{H}_{N})$ with the blocks placed along the diagonal.
@@
where $\hat{\mathbf{z}}_{t}$ is the current estimate of target position, $\mathbf{P}_{t}$ is the current estimation covariance matrix, and $\mathbf{A}$ is the process matrix of target motions. Then the update step of EKF further updates the covariance matrix using the $\mathbf{R}$ and $\mathbf{H}$ we have derived above:

### v3

@@ Abstract
Multi-robot target tracking finds extensive applications in different scenarios, such as environmental surveillance and wildfire management, which require the robustness of the practical deployment of multi-robot systems in uncertain and dangerous environments. Traditional approaches often focus on the performance of tracking accuracy with no modeling and assumption of the environments, neglecting potential environmental hazards which result in system failures in real-world deployments. To address this challenge, we investigate multi-robot target tracking in the adversarial environment considering sensing and communication attacks with uncertainty. We design specific strategies to avoid different danger zones and propose a multi-agent tracking framework under the perilous environment. We approximate the probabilistic constraints and formulate practical optimization strategies to address computational challenges efficiently. We evaluate the performance of our proposed methods in simulations to demonstrate the ability of robots to adjust their risk-aware behaviors under different levels of environmental uncertainty and risk confidence. The proposed method is further validated via real-world robot experiments where a team of drones successfully track dynamic ground robots while being risk-aware of the sensing and/or communication danger zones.
@@ 1 Introduction
Deployment of multiple robots to track real targets in practical scenarios remains challenging, as the environment or targets themselves can be dangerous or adversarial. Robots may be subject to attacks resulting in sensor damage, communication interruption, or other system faults. Therefore, tracking performance is no longer the only factor to consider, as the robots must also pay special attention to secure themselves from attacks. [9, 10, 11, 12, 13] have closely investigated multi-robot target tracking in the adversarial setting. [10] designed a distributed strategy based on divide-and-conquer that can protect robots from denial-of-service (DoS) attacks. However, it considers attacks in the worst case and overestimates the number of attacks. [10] then improved the overly conservative strategy with an algorithm that can infer the number of DoS attacks using information from 3-hop neighbors. [9] similarly focused on worst-case attacks but showed that the algorithm still gives superior performance even when robots are under non-worst-case attacks (such as random attacks).
Though awareness of risk has been investigated in the context of multi-robot systems in these prior works, the worst-case assumption is often adopted, leading to excessively conservative decision-making. In this paper, we alternatively consider risk in a probabilistic sense and model the safety requirement for robots, i.e., the risk level must not exceed a threshold, using chance-based constraints. Our problem is relevant to motion/trajectory planning under uncertainty [14, 15, 16, 17, 18][14] models an autonomous vehicle as a linear Gaussian system and imposes safety constraints for it by formulating a disjunctive convex optimization. [15] presented a probabilistic collision avoidance method to navigate micro aerial vehicles through dynamic obstacles. Linear approximation of the probabilistic collision constraints is applied such that the algorithm runs in real time. Inspired by these works, we apply the approximation techniques to alleviate the computational hardness of evaluating the risk constraints. Our contributions can be summarized as :
@@ 2.2 Sensing danger zone
where pdf is the probability density function. We have $\mathbf{w}_{i,\mathcal{S}_{l}}=\mathbf{x}_{\mathcal{S}_{l}}-\mathbf{x}_{i}$, and $\mathbf{w}_{i,\mathcal{S}_{l}}$ also follows a Gaussian distribution $\mathcal{N}(\mathbf{\mu}_{\mathcal{S}_{l}}~{}-~{}\mathbf{x}_{i},~{}\mathbf{\Sigma}_{\mathcal{S}_{l}})$. The probability is an integral of a multivariate Gaussian random variable over a disk-shaped region $\{\mathbf{w}|\lVert\mathbf{w}_{i,\mathcal{S}_{l}}\rVert\leq r_{l}\}$, which lacks a closed form solution and thus cannot be directly computed.
@@ 2.3 Communication danger zone
Figure 1: (a) Geometrical illustration for a communication danger zone where the position of the central attacking source is non-deterministic. The uncertainty is reflected as a blue radial gradient; (b) Shows the original integration region corresponding to the chance constraints; (c) Linearizes the integration region to compute an upper-bound of the multivariate integral. The line $\mathbf{a}^{\top}\mathbf{x}=b$ is tangent to the circle.
Next, we show the geometric conditions under which the jamming attack will be successful, given that the exact positions of the jamming sources are non-deterministic. We use robot $i$, its teammate $j$ for $j\neq i$, and the jamming source in zone $\mathcal{C}_{k}$ as an example, but the idea applies to all triplets. As shown in Fig. 1, when robot $i$ is inside a communication danger zone $\mathcal{C}_{k}$, the jamming source $k$ can attack the communication channel between robot $i$ and its teammate $j$. Let the distance between robot $i$ and $j$ be $c$, and the distance between robot $i$ and the jamming source $k$ be $a$. We encode the degree to which robot $i$ and $j$ can maintain reliable communication using the ratio of distances [19, 20], defined as:
@@ 2.4 Target Tracking Optimization with Danger Zones
With two types of danger zones introduced, we now present a nonlinear optimization program in Eq. 8 to model online target tracking under the threats of sensing and communication attacks. Eq. 8 optimizes tracking performance in the objective function and incorporates safety constraints originating from the danger zones. The optimization over control inputs $\mathbf{u}_{t}=[\mathbf{u}_{1,t},\dots,\mathbf{u}_{M,t}]^{\intercal}$ for the whole robot team is solved as follows,
@@
The objective function balances between minimizing tracking error (first term) and control efforts (second term) with constant weights $w_{1},w_{2}$. $f(\cdot,\cdot)$ represents the uncertainty in target state estimation. It is a nonlinear function dependent on robot positions at time $t+1$ and predicted target positions at time $t+1$, i.e. $\mathbf{x}_{t+1}$ and $\hat{\mathbf{z}}_{t+1}$. Eq. 8b describes the dynamics constraint of all the robots. $\mathbf{x}_{t+1}=[\mathbf{x}_{1,t+1},\dots,\mathbf{x}_{M,t+1}]^{\intercal}$ is the state of all robots at step $t+1$. We adopt a simple linear dynamics model, with process matrix $\mathbf{\Phi}$ and control matrix $\mathbf{\Lambda}$, but switching to more complex ones is straightforward. Eq. 8c is the constraint preventing robots from entering sensing danger zones. Eq. 8d is the constraint to preserve effective inter-robot communication in communication danger zones. We use $a_{ik}$ to denote the distance between robot $i$ and the jamming source of communication danger zone $\mathcal{C}_{k}$, and $c_{i}^{*}$ to denote the $c^{*}$ value of robot $i$ (see Eq. 6), both calculated using robot $i$’s position at time step $t+1$, i.e., $\mathbf{x}_{i,t+1}$. This problem is challenging to solve directly with probabilistic constraints (8c)(8d) and the objective function involved the evaluation towards robots’ position $\mathbf{x}_{t+1}$ and predicted positions of targets $\hat{\mathbf{z}}_{t+1}$. In the following section, we will discuss the reformulation and approximation of the optimization problem.
@@ 3.1 Chanced-based Constraints Approximation
$\text{Prob}(\mathbf{a}^{\top}\mathbf{x}\leq b)\leq\delta\iff\mathbf{a}^{\top}\mathbf{\mu}-b\geq\gamma,$ (9)
where $\gamma=\text{erf}^{-1}(1-2\delta)\sqrt{2\mathbf{a}^{\top}\mathbf{\Sigma}\mathbf{a}}$, $\text{erf}(\cdot)$ is the standard error function.
The chance constraints in Eq. 8c and Eq. 8d can be converted into deterministic constraints which consequently simplify the computation process. The integration areas for both types of danger zones are disk-shaped, as illustrated in Fig. 1. We employ a linear approximation of the integration area to calculate a conservative upper bound of the probability. This approximation is demonstrated in Fig. 1, where the integration boundary is approximated by a line $\mathbf{a}^{\top}\mathbf{x}=b$ tangent to the disk.
@@
In other words, it requires the approximate probability to be bounded by $\epsilon_{1}$. $\mathbf{a}_{i,\mathcal{S}_{l}}=(\mathbf{\mu}_{\mathcal{S}_{l}}-\mathbf{x}_{i})/\lVert\mathbf{\mu}_{\mathcal{S}_{l}}-\mathbf{x}_{i}\rVert$ is a unit vector. The chance constraint in Eq. 10 has the same form as Eq. 9, and thus is equivalent to a deterministic constraint,
@@ 3.2 Mission Objective
The objective function Eq. 8a involves minimizing the uncertainty of target state estimation when robots are at resultant position $\mathbf{x}_{t+1}$. We let $f(\cdot,\cdot)$ be the trace of the estimation covariance matrix for target positions calculated using $\mathbf{x}_{t+1}$ and $\hat{\mathbf{z}}_{t+1}$ with the latter being the estimated positions of targets at time $t+1$. In other words, the first term in Eq. 8a tries to find robots’ next-step positions that minimize the tracking error utilizing the team’s predicted information of the targets. In the upcoming part, we show that the trace function is non-convex in optimization variables.
To estimate the targets’ positions recursively, we assume each robot is equipped with one range and one bearing sensor, whose measurements are nonlinear in target positions. We use $\mathbf{z}_{t}$ to denote the positions of targets, with the position of a particular target $j$ being $\mathbf{z}_{j,t}=[x_{T_{j,t}},~{}y_{T_{j,t}}]^{\top}$, and use $\mathbf{x}_{i,t}=[x_{i,t},~{}y_{i,t}]^{\top}$ to denote the position of robot $i$. Based on the derivation in [21], robot $i$’s range and bearing measurements of target $j$ are $d_{ij}=\sqrt{\Delta x_{ij}^{2}+\Delta y_{ij}^{2}}$ and $\theta_{ij}=\text{arctan}(\frac{\Delta y_{ij}}{\Delta x_{ij}})-\phi_{i}$, where $\Delta x_{ij}=x_{T_{j}}-x_{i},{}\Delta y_{ij}=y_{T_{j}}-y_{i}$, and $\phi_{i}$ is the orientation of robot $i$. We drop the time step $t$ here for simplicity of notation. Robot $i$’s measurement of target $j$ is thus
@@
Here, $n_{d_{ij}}$ and $n_{\theta_{ij}}$ are noises in range and bearing measurements, respectively. We assume both follow zero-mean Gaussian distributions. Specifically, we let $n_{d_{ij}}\sim\mathcal{N}(0,R_{d,ij})$. The measurement noise covariance, $R_{d,ij}$, is a nonlinear function of the distance between robot $i$ and target $j$. We define it through its inverse as $R^{-1}_{d,ij}=a_{d}\cdot\text{exp}(-\lambda_{d}\lVert\mathbf{x}_{i}-\mathbf{z}_{j}\rVert)$, where $a_{d}$ and $\lambda_{d}$ are constant parameters determining the characteristic of noise. The bearing measurement noise covariance, $R_{\theta,ij}$, is defined similarly. The team-level measurement noise covariance matrix with respect to target $j$, i.e., $\mathbf{R}_{j}$, is thereby constructed as $\mathbf{R}_{j}^{-1}=\text{diag}(R^{-1}_{d,1j},R^{-1}_{\theta,1j},R^{-1}_{d,2j},R^{-1}_{\theta,2j},\cdots,R^{-1}_{d,Mj},R^{-1}_{\theta,Mj})$, and it follows that $\mathbf{R}=\text{diag}(R_{1},R_{2},\cdots,R_{N})$ represents the noise covariance matrix of the whole robot team measuring all targets.
To compute the uncertainty of target position estimation, we need measurement matrix for a combination of range and bearing sensors. By [21], we derive robot $i$’s measurement matrix of target $j$ as
@@
where $\mathbf{h}_{d_{ij}}$ and $\mathbf{h}_{\theta_{ij}}$ are column vectors corresponding to the range and bearing measurements. With $\mathbf{p}_{ij}=\mathbf{x}_{i}-\mathbf{z}_{j}$, the two measurement vectors are
@@
$\mathbf{J}$ is the $2\times 2$ rotational matrix applied at $-\frac{\pi}{2}$. We can concatenate the measurement matrices of target $j$ from each robot vertically to obtain the whole team’s measurement matrix as $\mathbf{H}_{j}=[\mathbf{H}_{1j};\mathbf{H}_{2j};\cdots;\mathbf{H}_{Mj}]$. The robot team’s measurement matrix of all targets are thereby constructed as $\mathbf{H}=\text{diag}(\mathbf{H}_{1},\mathbf{H}_{2},\cdots,\mathbf{H}_{N})$.
@@
where $\hat{\mathbf{z}}_{t}$ is the current estimate of target positions, $\mathbf{P}_{t}$ is the current estimation covariance matrix, and $\mathbf{A}$ is the process matrix of target motions. Then, the update step of EKF further updates the covariance matrix using $\mathbf{R}$ and $\mathbf{H}$ we have derived above: